OpenStudy (anonymous):

A rectangular athletic field is twice as long as it is wide. If the perimeter of the athletic field is 120 yards, what are the dimensions ? what is the width ?

OpenStudy (bibby):

What is the perimeter of a normal rectangle where

OpenStudy (e.cociuba):

Use "L" for length, "W" for width, and "P" for perimeter.

OpenStudy (e.cociuba):

P = 2L + 2W (perimeter is twice the length plus twice the width) L = 2W (twice as long as it is wide) P = 120 (the perimeter of the athletic field is 120 yards) 120 = 2L + 2W ------------------ Substitute 2W (from the second equation) for L: 120 = 2(2W) + 2W ------------------ Simplify: 120 = 4W + 2W 120=6W W = 120 / 6 = 20 The field is 20 yards wide. Since the length is twice the width, it's 40 yards long.